OpenStudy (goformit100):

A towel, when bleached, was found to have lost 20% of its length and 10% of its breadth. The percentage of decrease in area is: A. 10% B. 10.08% C. 20% D. 28%

OpenStudy (goformit100):

@Mimi_x3 @mayankdevnani

OpenStudy (mayankdevnani):

Let original length = x and original breadth = y.

OpenStudy (goformit100):

Thanks

OpenStudy (mayankdevnani):

Decrease in area = \[xy - [\frac{80}{100}x \times \frac{90}{100}y]\] \[=xy-\frac{18}{25}xy\] \[=\frac{7}{25}xy\] so , decrese %=\[\frac{7}{25}xy \times \frac{1}{xy} \times 100=28 % \]